OpenStudy (anonymous):

Which value of x makes this equation true? –8 + x = 0 A. 0 B. –8 C. No value of x will make it true. D. 8

OpenStudy (hoblos):

–8 + x = 0 add 8 for both sides, what will you get ?

OpenStudy (anonymous):

0

OpenStudy (hoblos):

8 - 8 + x = 0 +8 x = ?